madison started a bank account with $200. each year, she earns 5% in interest, which means in the account is multiplied by 1.05 each year.
a) what are the first three terms of this sequence?
b) write an equation to represent this sequence.
c) if madison does not withdraw money from her account, how much will she have in 10 years?

Answer by addingup(3677)

a) use the formula in b) to calculate 3 years
b) Amount = Principal(1+rt), where r is the interest rate and t is the time in years
c) 200(1+0.05(10)) = 200(1+0.5) = 200 x 1.5 = 300